# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v2

@@ 3.2 Utility Analysis of DPSGD
(Privacy-utility trade-off for DPSGD-DC). Assuming that the conditions in Theorem 3.9 are satisfied, for $\boldsymbol{\theta}_{T}$ output by DPSGD-DC with $L$-smooth and $\mu$-strongly convex losses, we have the following results.
  • For $\sigma_{\textup{DP}}^{2}=\mathcal{O}\bigl{(}\frac{\alpha}{\varepsilon}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{nb})\bigr{)}$, we have the following inequality
    $\leq\mathcal{O}\biggl{(}\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{D}{\sqrt{\eta T}}+\min\bigl{(}\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\textup{SGD}},\sqrt{\frac{\sigma_{\textup{SGD}}^{3}}{\mu C}}\bigr{)}$
    $+\frac{\sqrt{\eta}\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{d\alpha\eta\sqrt{L}}{\varepsilon C}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{nb})$
@@
    $+\sqrt{\frac{d\alpha\eta}{\varepsilon}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{nb})}\biggr{)}.$
  • If we further assume $b\leq\frac{n}{5}$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, and let $\sigma_{\textup{DP}}^{2}=\mathcal{O}\bigl{(}\frac{\alpha}{\varepsilon}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{n^{2}})\bigr{)}$, then we have the following inequality
    $\leq\mathcal{O}\biggl{(}\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{D}{\sqrt{\eta T}}+\min\bigl{(}\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\textup{SGD}},\sqrt{\frac{\sigma_{\textup{SGD}}^{3}}{\mu C}}\bigr{)}$
    $+\frac{\sqrt{\eta}\sigma_{\textup{SGD}}}{\sqrt{b}}+\frac{d\alpha\eta\sqrt{L}}{\varepsilon C}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{n^{2}})$
    $+\sqrt{\frac{d\alpha\eta}{\varepsilon}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{n^{2}})}\biggr{)}.$
Figure 1: The evolution of the: (a) privacy level and (b) training loss during DPSGD-GC with different batch sizes. The shaded error bars correspond to intervals covering $95$% of the realized values, obtained from the $10$ Monte Carlo trials. Note that the privacy and utility bounds in terms of the number of epochs, $E$, can be derived by substituting $T=\lceil\frac{n}{b}\rceil E$ into our main results.
@@
Corollary 3.12 suggests that the utility bound for DPSGD-DC comprises six terms. The first two terms capture optimization-related factors, reflecting the influence of clipping and projection on convergence behavior. The third term accounts for the inherent bias introduced by gradient clipping, while the fourth term reflects the stochastic noise resulting from SGD sampling. Finally, the last two terms quantify the impact of the injected DP noise. Unlike previous works on DPSGD-DC (Altschuler & Talwar, 2022; Altschuler et al., 2024; Kong & Ribero, 2024; Chien & Li, 2024), which primarily focused on RDP analysis without considering convergence performance, our results explicitly demonstrate how clipping and projection factors affect the utility of DPSGD-DC.
Note that the utility bounds can also be expressed in terms of the standard DP parameter $(\epsilon,\delta)$ by applying the conversion in Lemma 2.4.
@@
From Theorem 3.4, we establish a constant upper bound on privacy for DPSGD-DC based on the bounded domain diameter $D$, where smaller value of $D$ yields tighter privacy guarantees. Meanwhile, Corollary 3.12 demonstrates that when the conditions in Theorem 3.9 are satisfied, a smaller $D$ also leads to a lower upper bound on the convergence. It thus follows that reducing $D$ can enhance the utility guarantee in terms of the privacy-convergence trade-off under certain conditions.
@@ 4 Experiment Evaluation
In this section, we present empirical results estimating the privacy level via MIA. Following Kairouz et al. (2015), we estimate $(\epsilon,\delta)$-DP using the false positive rate (FPR) and false negative rate (FNR) of its attack model on the test data, applying the following formula:
@@
We train a private ResNet-18 network (He et al., 2016) as the target model using the Opacus library (Yousefpour et al., 2021) on the standard CIFAR10 dataset (Krizhevsky & Hinton, 2009). According to standard MIA protocols (Shokri et al., 2017), the training and test sets for each target and shadow model are randomly selected, equal in size, and mutually disjoint. We set the training set size to $10,000$. While the target model’s dataset does not overlap with those of the shadow models, different shadow models may partially share the same data. The shadow models are trained using the same architectures as the target model. For the attack model, we employ a two-layer multilayer perceptron (MLP) with $50$ hidden nodes and ReLU as activation functions.
We emphasize that the MIA serves primarily as a tool to provide a lower bound on privacy, capturing the trend with privacy level changes and validating the consistency of the theoretical bounds. By comparing experimental results with the theoretical trends, we aim for demonstrating the reasonableness of the derived privacy bounds, rather than offering an exact measure of privacy leakage. This approach allows us to examine how privacy evolves with varying experimental conditions under the same privacy attack.
Figure 2: The evolution of the privacy level during DPSGD-DC with different diameters of bounded domain. The red and blue lines correspond to the cases with batch sizes of $100$ and $400$, respectively.
Implementation details. For the target model, we use the SGD optimizer with the learning rate $\eta=0.1$, the noise level $\sigma_{\text{DP}}=0.002$, the clipping threshold $C=20$, and the confidence level $\delta=10^{-5}$. We train $10$ shadow models to simulate the behavior of the target model. For the attack classifier, we use the SGD optimizer with the initial learning rate of $0.01$, the weight decay of $5\times 10^{-4}$, and the momentum of $0.9$. The mini-batch size during training is set to be $100$. We predict the labels by selecting from the model’s output the class with the highest probability. Experimental results are reported by averaging over $10$ Monte Carlo trials. Our experiments are implemented in the PyTorch (Paszke et al., 2017) framework.
4.1 Effect of the Batch Size
We conduct experiments on DPSGD-GC with various batch sizes $b\in\{100,200,300,400,600,1000\}$. Figure 1 illustrates the evolution of the estimated privacy level and training loss per epoch. As expected, the estimated privacy level $\epsilon$ increases with the number of epochs, and larger batch sizes provide stronger privacy protection. These observations align with our theoretical results in Theorem 3.3. Additionally, DPSGD-GC converges more slowly with larger batch sizes, consistent with Lemma 3.7, further highlighting the trade-off between privacy and convergence.
4.2 Effect of the Bounded Domain Diameter
We then conduct experiments on DPSGD-DC using various diameters for the bounded domain $D\in\{20,60,100\}$ with batch sizes of $100$ and $400$, respectively. The estimated privacy parameter of DPSGD-DC for different bounded domain diameters $D$ is shown in Figure 2. As can be observed, DPSGD-DC provides stronger privacy guarantees with a smaller $D$, as limiting the domain diameter restricts the range of parameter variations to a narrower interval. Additionally, privacy leakage tends to stabilize as the number of epochs increases. This observation is not surprising, as the bounded domain assumption provides a constant upper privacy bound for DPSGD-DC, as demonstrated in Theorem 3.4.
@@ 5 Conclusions
In this paper, we have rigorously analyzed the privacy and utility guarantee of DPSGD, considering both gradient clipping (DPSGD-GC) and double clipping (DPSGD-DC). Our analysis extends the existing privacy bounds of DPSGD-GC and DPSGD-DC to general smooth and non-convex problems without relying on other assumptions. While previous works have focused solely on privacy characterization, we have also derived utility bounds corresponding to our RDP guarantees. This dual characterization admits a more comprehensive understanding of the privacy-utility trade-offs in DPSGD, providing valuable insights for developing more effective differentially private optimization algorithms.
Our work contributes towards the role that a deeper understanding of parameter projection and gradient clipping play in DPSGD-based algorithms. This has implications for both current applications and future developments in differentially private optimization.
Impact Statement
The goal of this paper is to advance the understanding of the privacy-utility trade-offs in DPSGD. As DPSGD is an essential component of introducing differential privacy in machine learning, our work strengthens data privacy protection by establishing theoretically guaranteed bounds.
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Preliminaries
A.1 Standard Facts
@@
Lemma A.1.
(Markov inequality). If $x$ is a non-negative random variable and $a>0$, then we have
@@
Lemma A.2.
(Projection operator is non-expansive). For any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$, we have
@@
A.2 Lemmas for Privacy Analysis
We then provide some supporting lemmas which are useful for our proof. First, the following lemma provides equivalent characterizations of the $\infty$-Wasserstein distance:
Definition A.6.
($\infty$-Wasserstein distance (Villani et al., 2009)). The $\infty$-Wasserstein distance between two distributions $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$ is defined as
@@ A.2 Lemmas for Privacy Analysis
where $(\boldsymbol{x},\boldsymbol{y})\sim\gamma$ indicates that the essential supremum is taken with respect to the joint distribution $\gamma$, and $\Gamma(\mathbb{P}_{\boldsymbol{\mu}},\mathbb{P}_{\boldsymbol{\nu}})$ represents the collection of the joint distribution with $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$ as marginals.
Lemma A.7.
(Equivalent definitions of $\infty$-Wasserstein distance (Feldman et al., 2018)). The following are equivalent for any distributions $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$:
@@
Lemma A.8.
(Post-processing inequality (Van Erven & Harremos, 2014)). Let $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ be two random variables. Then, for any (possibly random) function $\psi$ and $\alpha\geq 0$, we have
@@
Lemma A.9.
(Partial convexity in the second argument (Van Erven & Harremos, 2014)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}_{1}$, and $\boldsymbol{\nu}_{2}$ be three random variables. Then, for any $\alpha\geq 0$, the Rényi divergence is convex in its second argument, that is, the following inequality holds
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||(1-\lambda)\mathbb{P}_{\boldsymbol{\nu}_{1}}+\lambda\mathbb{P}_{\boldsymbol{\nu}_{2}})\leq(1-\lambda)\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}_{1}})+\lambda\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}_{2}}),$ (32)
@@
Lemma A.10.
(Strong composition for Rényi divergence (Mironov, 2017)). Let $\mathcal{M}=\left[\mathcal{M}_{1}(\mathcal{D}),\mathcal{M}_{2}\left(\mathcal{D}\right)\right]$ be a sequence of two randomized mechanisms. Then, for any $\alpha>1$, we have
$\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}(\mathcal{D})}||\mathbb{P}_{\mathcal{M}(\mathcal{D}^{\prime})})\leq\sup_{\boldsymbol{v}}\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}_{2}(\mathcal{D})|\mathcal{M}_{1}(\mathcal{D})=\boldsymbol{v}}||\mathbb{P}_{\mathcal{M}_{2}(\mathcal{D}^{\prime})|\mathcal{M}_{1}(\mathcal{D}^{\prime})=\boldsymbol{v}})+\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}_{1}(\mathcal{D})}||\mathbb{P}_{\mathcal{M}_{1}(\mathcal{D}^{\prime})}),$ (33)
@@
Definition A.11.
(Sampled Gaussian mechanism). Let $f:\mathcal{D}\rightarrow\mathbb{R}^{d}$ and $\mathcal{S}$ be a sample from $[n]$ where each $i\in[n]$ is chosen independently with probability $0<q\leq 1$ and $n=|\mathcal{D}|$. The SGM, parameterized with the noise scale $\sigma>0$, is defined as
@@
$\mathcal{D}_{\alpha}^{\text{SGM}}(q,\sigma)=\mathcal{D}_{\alpha}\left(\mathcal{N}(0,\sigma^{2})||(1-q)\mathcal{N}(0,\sigma^{2})+q\mathcal{N}(1,\sigma^{2})\right).$ (35)
@@
Lemma A.12.
(Upper bound on Rényi divergence of the SGM). If $q\leq 1/5$, $\sigma>4$, and $\alpha\leq\alpha^{*}(q,\sigma)$, i.e.,
@@
$\alpha$ $\leq\left(K^{2}\sigma^{2}/2-\log 5-2\log\sigma\right)/\left(K+\log(q\alpha)+1/(2\sigma^{2})\right),$
@@
A.3 Lemmas for Utility Analysis

### v3

@@ 3.2 Utility Analysis of DPSGD
(Privacy-utility trade-off for DPSGD-DC). Assuming that the conditions in Theorem 3.10 are satisfied, for DPSGD-DC with $L$-smooth and $\mu$-strongly convex population risk, $\sigma_{\textup{DP}}^{2}=\mathcal{O}\bigl(\frac{\alpha C^{2}}{\varepsilon nb}\min\{T,\overline{T}\}\bigr)$ and $\overline{T}=\frac{(1+\eta L)^{2}nbD^{2}}{\eta^{2}C^{2}}$, we have the following results:
$\min_{t\in[0,T]}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]\leq\mathcal{O}\biggl(\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{D}{\sqrt{\eta T}}$ (18)
$+\min\left\{\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\textup{SGD}},\sqrt{\frac{\sigma_{\textup{SGD}}^{3}}{\mu C}}\right\}+\frac{\sqrt{\eta}\sigma_{\textup{SGD}}}{\sqrt{b}}$
@@
$+\frac{\alpha d\eta C\sqrt{L}}{\varepsilon nb}\min\bigl\{T,\overline{T}\bigr\}+\sqrt{\frac{\alpha d\eta}{\varepsilon nb}\min\bigl\{T,\overline{T}\bigr\}}C\biggr).$
@@
Proposition 3.12 suggests that the utility bound for DPSGD-DC comprises six terms. The first two terms capture optimization-related factors, reflecting the influence of clipping and projection in convergence behavior. The third term accounts for the inherent bias introduced by gradient clipping, while the fourth term reflects the effect due to SGD sampling. Finally, the last two terms quantify the impact of the injected DP noise.
Following Proposition 3.12, we can derive the achievable privacy-utility trade-off for DPSGD-DC as
$\mathcal{O}\biggl(\max\biggl\{\frac{D^{2}dL\log(1/\delta)}{\epsilon^{2}n^{2}}$ $,\frac{\sigma_{\text{SGD}}^{3/2}D^{1/2}}{\mu^{1/2}}[\frac{dL\log(1/\delta)}{\epsilon^{2}n^{2}}]^{1/4}$ (19)
$,\frac{\sigma_{\text{SGD}}D\sqrt{d\log(1/\delta)}}{\sqrt{b}\epsilon}\biggr\}\biggr),$
and we provide the details in Appendix C.3.
@@
Unlike previous works on DPSGD-DC (altschuler2022privacy; altschuler2024privacy; kong2024privacy; chien2024convergent), which mainly focused on RDP analysis, our results explicitly demonstrate how gradient clipping and projection affect the utility of DPSGD-DC. As seen from Theorem 3.4, we establish an eventually constant upper bound on privacy loss dependent on the bounded domain diameter $D$, which yields tighter privacy guarantees with smaller $D$. Meanwhile, Proposition 3.12 demonstrates that when the conditions in Theorem 3.10 are satisfied, a smaller $D$ also leads to a lower upper bound on the utility, thus enhancing the privacy-utility trade-off. [^4] A brief discussion on selecting $D$ is provided in Appendix E.3.
Figure 2: The evolution of the: (a) estimated and (b) normalized theoretical privacy level during DPSGD-GC with different batch sizes. The shaded error bars correspond to intervals covering 9595% of the realized values, obtained from the $10$ Monte Carlo trials. Note that the privacy bounds in terms of the number of epochs, $E$, can be derived by substituting $T=\lceil\frac{n}{b}\rceil E$ into our main results.
Figure 3: The evolution of the privacy level during DPSGD-DC with different diameters of the bounded domain: (a) estimated by MIA and (b) normalized theoretical privacy level $\varepsilon$ with $\alpha=1.1$. The red and blue lines correspond to the cases with batch sizes of $100$ and $400$, respectively.
@@ 4 Experiment Evaluation
In this section, we present empirical results estimating the privacy level via MIA. Following kairouz2015composition, we estimate $(\epsilon,\delta)$-DP using the false positive rate (FPR) and false negative rate (FNR) of its attack model on the test data, applying the following formula:
@@
We emphasize that the MIA serves primarily as a tool to provide a lower bound on privacy, capturing the trend with privacy level changes and validating the consistency of the theoretical bounds. By comparing experimental results with the theoretical trends, we aim for demonstrating the reasonableness of the derived privacy bounds, rather than offering an exact measure of privacy leakage. This approach allows us to examine how privacy evolves with varying experimental conditions under the same privacy attack. Additional information on the MIA setting and implementation details can be found in Appendix D.
Effect of the Batch Size.
We evaluate DPSGD-GC with various batch sizes $b\in\{100,200,300,400,600,1000\}$. Figure 2 illustrates the evolution of the estimated and theoretical privacy level. We also report the evolution of training loss in Figure 4 in Appendix D.3. As expected, the estimated privacy level $\hat{\epsilon}$ increases with the number of epochs, and larger batch sizes provide stronger privacy protection. These observations align with our theoretical results in Theorem 3.3. Additionally, DPSGD-GC converges more slowly with larger batch sizes, consistent with Lemma 3.7, further highlighting the trade-off between privacy and utility.
Effect of the Bounded Domain Diameter.
We then conduct experiments on DPSGD-DC using various diameters for the bounded domain $D\in\{20,60,100\}$ with batch sizes of $100$ and $400$, respectively. The estimated and theoretical privacy parameter of DPSGD-DC for different bounded domain diameters $D$ is shown in Figure 3. As can be observed, DPSGD-DC provides stronger privacy guarantees with a smaller $D$, as limiting the domain diameter restricts the range of parameter variations to a narrower interval. Additionally, privacy leakage tends to stabilize as the number of epochs increases. This observation is not surprising, as the bounded domain assumption provides a constant upper privacy bound for DPSGD-DC, as shown in Theorem 3.4.
@@ 5 Conclusions
In this paper, we have rigorously analyzed the privacy and utility guarantee of DPSGD, considering both gradient clipping (DPSGD-GC) and double clipping (DPSGD-DC). Our analysis extends the existing privacy bounds of DPSGD to general smooth and non-convex problems without relying on other assumptions. While previous works have focused solely on privacy characterization, we have also derived utility bounds corresponding to our RDP guarantees. This dual characterization admits a more comprehensive understanding of the privacy-utility trade-offs in DPSGD, providing valuable insights for developing more effective differentially private optimization algorithms.
Acknowledgments
The authors would like to thank the anonymous reviewers, SPC, and AC for their insightful comments. The work of H. Xing was supported in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2025A1515010123, and in part by the Guangzhou Municipal Science and Technology Project under Grants 2024A04J4527 and 2023A03J0663.
@@ Appendix A Preliminaries
A.1 Privacy Amplification by Iteration
The privacy amplification by iteration framework relies on two key lemmas, focusing specifically on Gaussian noise, which suffices for our purposes. First, Lemma 2.6 stated in the main body will prove to be useful for analyzing shifted Rényi divergence between two distributions convolved with Gaussian noise.
Lemma A.1 (Contraction-reduction (feldman2018privacy)).
Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Suppose $\phi(\cdot)$, $\phi^{\prime}(\cdot)$ are two contractive functions and let $\sup_{\boldsymbol{x}}\|\phi(\boldsymbol{x})-\phi^{\prime}(\boldsymbol{x})\|\leq s$. Then, for any $z\geq 0$, we have
Altschuler et al. (altschuler2022privacy; altschuler2024privacy) provide a mild generalization of this lemma, extending it to be suitable for random contraction maps.
A.2 Standard Facts
@@
Lemma A.2 (Markov inequality).
If $x$ is a non-negative random variable and $a>0$, then we have
@@
Lemma A.3 (Projection operator is non-expansive).
For any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$, we have
@@
A.3 Lemmas for Privacy Analysis
We then provide some supporting lemmas that are useful for our proof. First, the following lemma provides equivalent characterizations of the $\infty$-Wasserstein distance:
Definition A.7 ($\infty$-Wasserstein distance (villani2009wasserstein)).
The $\infty$-Wasserstein distance between two distributions $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$ is defined as
@@ A.3 Lemmas for Privacy Analysis
where $(\boldsymbol{x},\boldsymbol{y})\sim\gamma$ indicates that $(\boldsymbol{x},\boldsymbol{y})$ follows the joint distribution $\gamma$, and $\Gamma(\mathbb{P}_{\boldsymbol{\mu}},\mathbb{P}_{\boldsymbol{\nu}})$ represents the collection of all joint distributions with $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$ being as marginals of each random variable.
Lemma A.8 (Equivalent definitions of $\infty$-Wasserstein distance (feldman2018privacy)).
The following are equivalent for any distributions $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$:
@@
Lemma A.9 (Post-processing inequality (van2014renyi)).
Let $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ be two random variables. Then, for any (possibly random) function $\psi$ and $\alpha\geq 0$, we have
@@
Lemma A.10 (Partial convexity in the second argument (van2014renyi)).
Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}_{1}$, and $\boldsymbol{\nu}_{2}$ be three random variables. Then, for any $\alpha\geq 0$, the Rényi divergence is convex in its second argument, that is, the following inequality holds
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||(1-\lambda)\mathbb{P}_{\boldsymbol{\nu}_{1}}+\lambda\mathbb{P}_{\boldsymbol{\nu}_{2}})$ (29)
$\leq(1-\lambda)\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}_{1}})+\lambda\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}_{2}}),$
@@
Lemma A.11 (Strong composition for Rényi divergence (mironov2017renyi)).
Let $\mathcal{M}=\left[\mathcal{M}_{1}(\mathcal{D}),\mathcal{M}_{2}\left(\mathcal{D}\right)\right]$ be a sequence of two randomized mechanisms. Then, for any $\alpha>1$, we have
$\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}(\mathcal{D})}||\mathbb{P}_{\mathcal{M}(\mathcal{D}^{\prime})})$ (30)
$\leq\sup_{\boldsymbol{v}}\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}_{2}(\mathcal{D})|\mathcal{M}_{1}(\mathcal{D})=\boldsymbol{v}}||\mathbb{P}_{\mathcal{M}_{2}(\mathcal{D}^{\prime})|\mathcal{M}_{1}(\mathcal{D}^{\prime})=\boldsymbol{v}})$
$+\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}_{1}(\mathcal{D})}||\mathbb{P}_{\mathcal{M}_{1}(\mathcal{D}^{\prime})}),$
@@
Definition A.12 (Sampled Gaussian mechanism).
Let $f(\cdot):\mathcal{D}\rightarrow\mathbb{R}^{d}$ and $\mathcal{S}$ be a sample from $[n]$ where each $i\in[n]$ is chosen independently with probability $0<q\leq 1$ and $n=|\mathcal{D}|$. The SGM, parameterized with the noise scale $\sigma>0$, is defined as
@@
$\mathcal{D}_{\alpha}^{\text{SGM}}(q,\sigma)$ (32)
$=\mathcal{D}_{\alpha}\left(\mathcal{N}(0,\sigma^{2})||(1-q)\mathcal{N}(0,\sigma^{2})+q\mathcal{N}(1,\sigma^{2})\right).$
@@
Lemma A.13 (Upper bound on Rényi divergence of the SGM (mironov2019r)).
If $q\leq 1/5$, $\sigma>4$, and $\alpha\leq\alpha^{*}(q,\sigma)$, i.e.,
@@
$\alpha$ $\leq\frac{\left(K^{2}\sigma^{2}/2-\log 5-2\log\sigma\right)}{\left(K+\log(q\alpha)+1/(2\sigma^{2})\right)},$
@@
A.4 Lemmas for Utility Analysis